A substance of mass 0.4kg having a density 2700kg/m3 occupies an area 0.02m2. Determine the thickness of the substance.
1
Expert's answer
2020-06-08T10:25:36-0400

Given data

Density of substance is "\rho=2700 kg/m^3"

Area is "A=0.02 m^2"

The expression for the density of the substance is

"\rho=\frac{m}{At}"

So the thickness of the substance is

"t=\frac{m}{\rho A}=\frac{0.4}{(2700)(0.02)}=7.41*10^{-3} m"